Wallace and Vernonville are 208 miles apart. A car leaves Wallace travailing towards Vernonville, and another car leaves Vernonville at the same time, traviling towards Wallace. The car leaving Wallace avarages 10 miles per hour more than the other, and they meet after 1 hour and 36 minutes. What are the avetage speeds of the cars?
Answer by mananth(16949)

car leaving wallace x mph
Car leaving Vernonville = x-10 mph

Towards each other so speed = x+x-10==>2x-10
D= 208
t=1h 36 m ===> 8/5 hours
speed = 2x-10
speed = d/t
speed = 208/(8/5)
2x-10 = 208*5/8
2x-10 =26*5
2x-10=180
2x=180+10
2x=190
x=95 speed of one car
other car speed = 85 mph